Physics 335—Thermal Physics Homework Assignment #10 Due November 20, 2024, 11:40 a.m.

Ch 7:

Problem 7.8, 7.9, and 7.15 (20 pts.xeach).

Note: This week’s assignment is short to allow you more time to work on your presentation problem.